A certain game consists of rolling a single fair die. If a four or five comes​ up, you win 6 dollars; otherwise, you lose 3 dollars. Find the expected winnings for this game. Is this a fair​ game?
Answer by math_helper(2461)


Expected winnings = (2/6)($6) - (4/6)($3) = 2-2 = 0
So yes, this is a fair game.